# Scaled Block Vecchia Approximation for High-Dimensional Gaussian Process Emulation on GPUs

### v2

@@ 1. Introduction
A common approach to building computer model emulators involves fitting a Gaussian Process (GP) to data generated by running the simulator at a predefined set of input values (design points), effectively learning a mapping from the model input space to the model output space (Andrianakis and Challenor, 2012). GP-based emulators enhance interpretability and facilitate efficient analysis, particularly when simulating scenarios across a broad parameter space. For example, variable importance (Linkletter et al., 2006; Saltelli et al., 2010) and sequential optimization (Winkel et al., 2021) may only be feasible using a cheap-to-evaluate GP emulator of the underlying expensive-to-evaluate simulator.
@@
Among these approximation methods, Classic Vecchia (CV) approximation demonstrates superior advantages (Hazra et al., 2024). In (Katzfuss et al., 2022), Katzfuss et al. introduced the Vecchia approximation for GP emulation in the context of computer experiments and proposed the Scaled Vecchia (SV) method to enhance approximation accuracy in high-dimensional settings. However, their study was limited to small-scale problems and relied entirely on CPU-based computations. In contrast, efforts have been made to accelerate the classic Vecchia algorithm on GPUs, as seen in (Pan et al., 2024), which subsequently introduced the Block Vecchia (BV) approach (Pan et al., 2025) to enhance computational speed further while preserving approximation accuracy. Nevertheless, the use of CV approximation for high-dimensional GP emulation on GPUs remains unexplored. This gap presents an opportunity to develop a scalable GPU-based method that can efficiently address the computational and memory challenges posed by high-dimensional inputs.
In this work, we propose a novel algorithm called the Scaled Block Vecchia (SBV) approximation for distributed GPU-based systems. SBV integrates the Block Vecchia (BV) algorithm from (Pan et al., 2025) with the Scaled Vecchia approach introduced in (Katzfuss et al., 2022), while leveraging GPU acceleration techniques from (Pan et al., 2024). To the best of our knowledge, this work represents the first distributed implementation of any variant of the Vecchia-based GP algorithm. Leveraging the Message Passing Interface (MPI), we parallelize computations across multiple GPU-enabled nodes, enabling scalable performance well beyond the capabilities demonstrated in prior studies.
To validate the accuracy and robustness of our algorithm, we perform GP simulations across a range of parameter settings and benchmark its performance on the satellite drag dataset, a standard high-dimensional testbed for evaluating GP-based models (Sun et al., 2019). As an application in computer model emulation, we apply our algorithm to MetaRVM, a compartmental simulation model designed for generic respiratory virus diseases (Fadikar et al., 2025). Experimental results demonstrate both high predictive accuracy and strong scalability, with successful runs on up to 512 A100/GH200 GPUs and problem sizes reaching 2.562.56B data points. Additionally, our approach achieves a noticeable reduction in power consumption compared to the state-of-the-art exact GP framework on GPUs, i.e., ExaGeoStat (Abdulah et al., 2018a).
@@ 2. Contributions
  • We propose the Scaled Block Vecchia (SBV) approximation, a novel GP algorithm that integrates anisotropic input scaling with block-based conditioning to reduce both computational and memory complexity. SBV improves approximation accuracy in high-dimensional GPs.
  • We develop the first distributed implementation of any Vecchia-based GP variant designed for modern GPU-accelerated HPC systems. Our implementation employs MPI for inter-node parallelism and the MAGMA library (Abdelfattah et al., 2021; Agullo et al., 2009) for fine-grained batched GPU-accelerated linear algebra operations.
@@
  • We use and assess SBV to emulate the MetaRVM epidemiological simulation using 5050M samples and 10-dimensional inputs, achieving high predictive accuracy and practical scalability.
  • We demonstrate both weak and strong scaling on the JURECA-DC GPU (AMD EPYC + A100) and JUPITER (GH200 Superchip) systems at Jülich Supercomputing Centre (JSC), with experiments scaling up to $512$ nodes and problem sizes reaching up to 2.562.56B data points.
@@ 3. Related Work
CV approximation: CV approximation  (Vecchia, 1988; Katzfuss and Guinness, 2021) reduces the computational complexity of GPs by factorizing the joint distribution into a product of conditional distributions. Recent advancements have enhanced the effectiveness of the CV approximation through improved data point ordering and neighbor selection (Guinness, 2018). The BV approximation (Pan et al., 2025) builds on this by conditioning on blocks of observations rather than individual points, enabling batched linear algebra operations on GPUs and improving throughput and memory efficiency in high-performance environments (Pan et al., 2024). More recently, the SV method was introduced to address challenges associated with high-dimensional inputs by applying anisotropic input scaling (Katzfuss et al., 2022), which enables the model to capture directional-varying correlations. Our work extends these developments by integrating block-based conditioning and anisotropic scaling into a unified framework, i.e., SBV, and presenting a distributed implementation optimized for multi-GPU systems.
@@ 4.1. Gaussian Processes (GPs)
where the conditional mean $\mu^{*}$ and covariance $\Sigma^{*}$ are derived as $\mu^{*}=\Sigma_{\bm{\theta},*}^{\top}\Sigma_{\bm{\theta}}^{-1}\bm{y},$ $\Sigma^{*}=\Sigma_{\bm{\theta}}^{*}-\Sigma_{\bm{\theta},*}^{\top}\Sigma_{\bm{\theta}}^{-1}\Sigma_{\bm{\theta},*}.$ The GPs have computational complexity $\mathcal{O}(n^{3})$ and memory complexity $\mathcal{O}(n^{2})$, making it infeasible on large scale problems.
@@ 4.3. Kullback-Leibler (KL) Divergence
The KL divergence measures how one probability distribution diverges from a second probability distribution (Cover and Thomas, 2006). For two Gaussian distributions $\mathcal{N}_{0}$ and $\mathcal{N}_{1}$ with zero mean and covariance matrices $\Sigma_{0}$ and $\Sigma_{1}$, the KL divergence is (Murphy, 2012):
$D_{KL}(\mathcal{N}_{0}\|\mathcal{N}_{1})=\frac{1}{2}\left\{\text{tr}(\Sigma_{1}^{-1}\Sigma_{0})-n+\log\frac{|\Sigma_{1}|}{|\Sigma_{0}|}\right\}.$
@@ 4.4. Scaled Kernel Function
In high-dimensional spaces, covariance functions often require anisotropic scaling, where each input dimension is scaled according to its relevance to the output (Katzfuss et al., 2022). This allows the model to reflect varying sensitivities across dimensions and improves both accuracy and efficiency. The scaled covariance function is defined as
(5) $K_{\bm{\theta}}(\bm{x}_{k},\bm{x}_{k^{\prime}})=f\left(\left(\sum_{i=1}^{d}\frac{(x_{ki}-x_{k^{\prime}i})^{2}}{\beta_{i}^{2}}\right)^{\frac{1}{2}}\right),$
@@ 5. Distributed Scaled Block Vecchia Algorithm
This section presents the proposed distributed Scaled Block Vecchia (SBV) algorithm, consisting of data preprocessing and batched GPU-based log-likelihood computation. We then analyze the computational complexity of the algorithm, highlighting efficient GPU utilization.
5.1. GP Modeling
The distributed SBV algorithm consists of four key components: scaling and partitioning, Random Anchor Clustering (RAC), filtered subset selection for NNS, and batched GPU computation for optimizing the log-likelihood. The preprocessing steps, i.e., scaling and partitioning, RAC, and filtered NNS, are performed only once and executed on the CPU to prepare the desired data structure. The resulting structured data is then transferred to the GPU, allowing for the efficient execution of hundreds of parameter optimization iterations without redundant data movement. The distributed SBV algorithm is detailed in Algorithm 1, and the pipeline is illustrated in Figure 2, where Steps (2), (3), and (4) are the Block Vecchia (BV) algorithm pipeline.
@@ 5.1. GP Modeling
1:Input: Data $\{\bm{x}_{i},y_{i}\}_{i=1}^{n}$, dimension $d$, total block count $K$, local clusers $k_{p}$, nearest neighbors $m_{\text{est}}$, workers $P$, covariance function $K_{\theta}$, scaling parameters $\{\beta_{i}\}_{i=1}^{d}$.
@@
5.1.1. Data Partitioning and Scaling
Each worker $p$ begins by loading its assigned data subset $\bm{X}_{p}$. Firstly, the points are partitioned into several segments along their most relevant dimension, as shown in line 4 of Algorithm 2. Here, the most relevant dimension is denoted by $d_{max}$, and MPI_Alltoall is used to redistribute the dataset $X_{1:P}$, where $x_{p,i,d_{max}}$ refers to the $d_{max}$-th component of the $i$-th data point on worker $p$. Secondly, input dimensions are rescaled using anisotropic parameters $\bm{\beta}=[\beta_{1},\ldots,\beta_{d}]$. This scaling transforms each dimension $x_{ij}:=x^{org}_{ij}/\beta_{j}$, ensuring that the relevance of different dimensions is appropriately reweighted (see line 7 in Algorithm 2).
Figure 2 gives an example of partitioning within the unit square $0,1]^{2}$. During data loading, each color represents the portion assigned to a specific worker; for example, the green points are read by Worker 1. In the scaling and partition, the input space is partitioned into four segments along the most relevant dimension, $\beta_{2}=0.3$. After that, the input space is rescaled from the unit square to the rectangle $(0,1.5)\times(0,0.3)$, indicating that the first dimension (width) is more relevant to the response than the second (height). This rescaling and partitioning strategy ensures balanced and uniform data distribution across workers in the algorithm, enhancing computational efficiency while preserving accuracy.
Algorithm 2 Partitioning and Scaling $\mathcal{PS}(X^{org}_{1:P})$
1:Input: Complete dataset $X^{org}_{1:P}$.
2:Output: Scaled and redistributed $X_{1:P}$. Relevant dimension $d_{max}=\arg\max_{i}\{1/\beta_{i},i=1,2,\dots,d\}.$
@@ 5.1.1. Data Partitioning and Scaling
7:  Dimension scaling, $X_{p,j}:=X^{org}_{p,j}/\beta_{j}$, $j=1,2,\ldots,d$.
@@
9:Return: Redistributed and scaled $X_{1:P}$
5.1.2. Random Anchor Clustering (RAC)
Next, small blocks $\bm{B}_{p}$ are created, defined as $\bm{B}_{p}:=\{\bm{B}_{p,1},\ldots,\bm{B}_{p,k_{p}}\}$, for each partition by RAC independently. The RAC method is detailed in Algorithm 3. Here, we replace the K-means clustering in block Vecchia GPs (Pan et al., 2025) with RAC to reduce the computational cost induced by K-means for extensive datasets while maintaining comparable approximation accuracy. The RAC algorithm begins by randomly selecting $k_{p}$ anchors (centers) for each worker $p=1,2,\dots,P$ in parallel. These anchors serve as the centers of blocks. Subsequently, each $x_{p,i}$ is assigned to the $j$th block based on their distance, as shown in line 5 of Algorithm 3, thus constructing clusters, $B_{p}:=(B_{p,1},B_{p,2},\ldots,B_{p,k_{p}})$ with $p=1,2,\dots,P$. The RAC method does not involve any communication between workers, considering that the size of clusters is typically much smaller than the overall dataset. For example, a single cluster may contain only 100, while the entire dataset consists of 10 million locations. In Step 2 of Figure 2, each worker creates 4 blocks in their data partition.
Algorithm 3 RAC algorithm $\mathcal{C}(X_{1:P})$
1:Input: $X_{1:P}$ full dataset.
2:Output: $B_{1:P}$ set of cluster for each node.

### v3

@@ 1. Introduction
A common approach to building computer model emulators is to fit a Gaussian Process (GP) to data generated by running the simulator at a predefined set of input values (design points), thereby learning a mapping from the model input space to the model output space (Andrianakis and Challenor, 2012). GP-based emulators enhance interpretability and facilitate efficient analysis, particularly when simulating scenarios across a broad parameter space. For example, variable importance (Saltelli et al., 2010) and sequential optimization (Winkel et al., 2021) may only be feasible using a cheap-to-evaluate GP emulator of the underlying expensive-to-evaluate simulator. While methods such as RBF-based models (Acosta, 1995; Arora et al., 2023) can offer flexible function approximation for such endeavors, they are fundamentally deterministic heuristics that lack marginal likelihood formulations, statistically grounded parameter estimation, and coherent uncertainty quantification. As a result, they are not well aligned with the probabilistic and inference-driven objectives of this work. In contrast, GP modeling provides a fully probabilistic framework with principled uncertainty quantification and likelihood-based inference, which are essential when studying scalable approaches to inference and made possible by covariance approximations such as Vecchia-based methods (Vecchia, 1988; Katzfuss and Guinness, 2021).
@@
Among these approximation methods, Classic Vecchia (CV) approximation demonstrates superior advantages. In (Katzfuss et al., 2022), Katzfuss et al. introduced the Vecchia approximation for GP emulation in the context of computer experiments and proposed the Scaled Vecchia (SV) method to enhance approximation accuracy in high-dimensional settings. However, their study was limited to small-scale problems and relied entirely on CPU-based computations. In contrast, efforts have been made to accelerate the classic Vecchia algorithm on GPUs, as in (Pan et al., 2024), which subsequently introduced the Block Vecchia (BV) approach (Pan et al., 2025) to further enhance computational speed while preserving approximation accuracy. Nevertheless, the use of CV approximation for high-dimensional GP emulation on GPUs remains unexplored. This gap presents an opportunity to develop a scalable GPU-based method that can efficiently address the computational and memory challenges posed by high-dimensional inputs.
In this work, we propose a novel algorithm, the Scaled Block Vecchia (SBV) approximation, for distributed GPU-based systems. SBV integrates the Block Vecchia (BV) algorithm from (Pan et al., 2025) with the Scaled Vecchia approach introduced in (Katzfuss et al., 2022), while leveraging GPU acceleration techniques from (Pan et al., 2024). To the best of our knowledge, this work represents the first distributed implementation of any variant of the Vecchia-based GP algorithm. Leveraging the Message Passing Interface (MPI), we parallelize computations across multiple GPU-enabled nodes, achieving scalable performance well beyond that demonstrated in prior studies.
To validate the accuracy and robustness of our algorithm, we perform GP simulations across a range of parameter settings and benchmark its performance on the satellite drag dataset, a standard high-dimensional testbed for evaluating GP-based models (Sun et al., 2019). As an application in computer model emulation, we apply our algorithm to MetaRVM, a compartmental simulation model designed for generic respiratory virus diseases (Fadikar et al., 2025). Experimental results demonstrate both high predictive accuracy and strong scalability, with successful runs on up to 512 A100/GH200 GPUs and problem sizes of up to 2.562.56B data points. Additionally, our approach achieves a noticeable reduction in power consumption compared to the state-of-the-art exact GP framework on GPUs, i.e., ExaGeoStat (Abdulah et al., 2018a).
@@ 2. Contributions
  • We propose the Scaled Block Vecchia (SBV) approximation, a novel GP algorithm that integrates anisotropic input scaling with block-based conditioning. While BV enables scalable inference for large spatial datasets in low-dimensional settings and SV improves approximation accuracy in high-dimensional inputs, the regime of high-dimensional and large-scale GP inference remains largely unresolved. SBV bridges this gap by combining the advantages of SV and BV, improving approximation accuracy while reducing both computational and memory complexity.
  • We develop efficient clustering and neighbor-selection algorithms, including Random Anchor Clustering (RAC) and a filtered block-wise KNN construction, that reduce the clustering and neighbor-search costs to approximately linear complexity. These algorithms make SBV practical for billion-point, high-dimensional datasets where standard clustering and nearest-neighbor construction would otherwise be computationally prohibitive.
  • We develop a distributed implementation of SBV for modern GPU-accelerated HPC systems. Our implementation employs MPI for inter-node parallelism and the MAGMA library (Abdelfattah et al., 2021; Agullo et al., 2009) for fine-grained batched GPU-accelerated linear algebra operations.
@@
  • We apply SBV to emulate the MetaRVM epidemiological simulation using 5050M samples and $10$-dimensional inputs, achieving high predictive accuracy and practical scalability.
  • We demonstrate both weak and strong scaling on the JURECA-DC GPU (AMD EPYC + A100) and JUPITER (GH200 Superchip) systems at Jülich Supercomputing Centre (JSC), with experiments scaling up to $512$ nodes and problem sizes reaching 2.562.56B data points.
@@ 3. Related Work
CV approximation: CV approximation  (Vecchia, 1988; Katzfuss and Guinness, 2021) reduces the computational complexity of GPs by factorizing the joint distribution into a product of conditional distributions. Recent advancements have enhanced the effectiveness of the CV approximation by improving data-point ordering and neighbor selection (Guinness, 2018). The BV approximation (Pan et al., 2025) builds on this by conditioning on blocks of observations rather than individual points, enabling batched linear algebra operations on GPUs and improving throughput and memory efficiency in high-performance environments (Pan et al., 2024). More recently, the SV method was introduced to address challenges associated with high-dimensional inputs by applying anisotropic input scaling (Katzfuss et al., 2022), which enables the model to capture directional-varying correlations. Our work extends these developments by integrating block-based conditioning and anisotropic scaling into a unified framework, i.e., SBV, and presenting a distributed implementation optimized for multi-GPU systems.
Vecchia-Based Applications and Extensions Recent research has increasingly adopted the Vecchia approximation as a scalable computational primitive within broader probabilistic modeling frameworks. For example, Vecchia approximations have been incorporated into deep Gaussian processes to enable fully Bayesian posterior inference with near-linear scaling for large computer experiments (Sauer et al., 2023). In the context of latent Gaussian process models with non-Gaussian likelihoods, iterative methods have been developed to accelerate Vecchia-Laplace approximations using conjugate gradients and stochastic trace estimation (Kündig and Sigrist, 2025). It has also been extended to scalable Gaussian process regression with variable selection through penalized likelihood formulations and mini-batch subsampling (Cao et al., 2022), and adapted to large-scale Bayesian optimization by integrating approximate nearest-neighbor search and variance recalibration strategies (Jimenez and Katzfuss, 2023). Beyond regression settings, Vecchia likelihood approximations have been employed for computationally tractable inference in high-dimensional spatial max-stable models (Huser et al., 2024), and have recently enabled linear-cost estimation of multivariate normal probabilities via sparse inverse Cholesky structures (Cao and Katzfuss, 2025). Previous studies have shown that the Vecchia approximation is a versatile and scalable tool for probabilistic modeling in various statistical fields. In contrast, our work pursues a different goal. We focus on likelihood-based Gaussian process emulation for high-dimensional inputs, combining scaled and block conditioning strategies, distributed-memory parallelization using MPI, GPU-accelerated batched dense linear algebra, and energy-efficient analysis on modern GPU supercomputers with billion-scale datasets. To our knowledge, this is the first distributed GPU implementation of a Vecchia-based Gaussian process method that scales to billions of data points, advancing Vecchia approximations from methodological scalability to extreme-scale HPC through accelerators.
@@ 4.1. Gaussian Processes (GPs)
where the conditional mean $\mu^{*}$ and covariance $\Sigma^{*}$ are derived as $\mu^{*}=\Sigma_{\bm{\theta},*}^{\top}\Sigma_{\bm{\theta}}^{-1}\bm{y},$ $\Sigma^{*}=\Sigma_{\bm{\theta}}^{*}-\Sigma_{\bm{\theta},*}^{\top}\Sigma_{\bm{\theta}}^{-1}\Sigma_{\bm{\theta},*}.$ The GPs have computational complexity $\mathcal{O}(n^{3})$ and memory complexity $\mathcal{O}(n^{2})$, making it infeasible on large scale problems (Williams and Rasmussen, 2006; Vecchia, 1988).
@@ 4.3. Kullback-Leibler (KL) Divergence
The KL divergence measures how one probability distribution diverges from a second probability distribution (Cover and Thomas, 2006). For two Gaussian distributions $\mathcal{N}_{0}$ and $\mathcal{N}_{1}$ with zero mean and covariance matrices $\bm{\Sigma}_{0}$ and $\bm{\Sigma}_{1}$, the KL divergence is (Murphy, 2012):
$D_{KL}(\mathcal{N}_{0}\|\mathcal{N}_{1})=\frac{1}{2}\left\{\text{tr}(\bm{\Sigma}_{1}^{-1}\bm{\Sigma}_{0})-n+\log\frac{|\bm{\Sigma}_{1}|}{|\bm{\Sigma}_{0}|}\right\}.$
@@ 4.4. Scaled Kernel Function
In high-dimensional spaces, covariance functions often require anisotropic scaling, where each input dimension is scaled according to its relevance to the output (Katzfuss et al., 2022). This allows the model to reflect varying sensitivities across dimensions, thereby improving both accuracy and efficiency. The scaled covariance function is defined as
(5) $K_{\bm{\theta}}(\bm{x}_{k},\bm{x}_{k^{\prime}})=f\left(\left(\sum_{i=1}^{d}\frac{(x_{ki}-x_{k^{\prime}i})^{2}}{\bm{\beta}_{i}^{2}}\right)^{\frac{1}{2}}\right),$
@@ 5. Distributed Scaled Block Vecchia Algorithm
This section presents the proposed distributed Scaled Block Vecchia (SBV) algorithm. The algorithm comprises four key components: scaling and partitioning, Random Anchor Clustering (RAC), filtered subset selection for NNS, and batched GPU computation to optimize the log-likelihood. The preprocessing steps, i.e., scaling and partitioning, RAC, and filtered NNS, are performed once on the CPU to prepare the desired data structure. The resulting structured data is then transferred to the GPU, enabling efficient execution of hundreds of parameter optimization iterations without redundant data movement. The distributed SBV algorithm is detailed in Algorithm 1, and the pipeline is illustrated in Figure 2, where Steps (2), (3), and (4) are the Block Vecchia (BV) algorithm pipeline.
@@
1:Input: Data $\{\bm{x}_{i},y_{i}\}_{i=1}^{n}$, dimension $d$, total block count $K$, nearest neighbors $m_{\text{est}}$, workers $P$, covariance function $K_{\theta}$, scaling parameters $\{\beta_{i}\}_{i=1}^{d}$.
@@
5.1. Data Partitioning and Scaling
Each worker $p$ begins by loading its assigned data subset $\bm{X}_{p}$. Firstly, the points are partitioned into several segments along their most relevant dimension, as shown in line 4 of Algorithm 2. Here, the most relevant dimension is denoted by $d_{max}$, and MPI_Alltoall is used to redistribute the dataset $\bm{X}_{1:P}$, where $x_{p,i,d_{max}}$ refers to the $d_{max}$-th component of the $i$-th data point on worker $p$. Secondly, input dimensions are rescaled using anisotropic parameters $\bm{\beta}=[\beta_{1},\ldots,\beta_{d}]$. This scaling transforms each dimension $x_{ij}:=x^{org}_{ij}/\beta_{j}$, ensuring that the relevance of different dimensions is appropriately reweighted (see line 7 in Algorithm 2).
Figure 2 gives an example of partitioning within the unit square $0,1]^{2}$. During data loading, each color denotes the portion assigned to a specific worker; for example, the green points are assigned to Worker 1. In the scaling and partitioning step, the input space is partitioned into four segments along the most relevant dimension, with $\beta_{2}=0.3$. After that, the input space is rescaled from the unit square to the rectangle $(0,1.5)\times(0,0.3)$, indicating that the first dimension (width) is more relevant to the response than the second (height). This rescaling and partitioning strategy ensures balanced and uniform data distribution across workers in the algorithm, enhancing computational efficiency while preserving accuracy.
Algorithm 2 Partitioning and Scaling $\mathcal{PS}(\bm{X}^{org}_{1:P})$
1:Input: Complete dataset $\bm{X}^{org}_{1:P}$.
2:Output: Scaled and redistributed $\bm{X}_{1:P}$. Relevant dimension $d_{max}=\arg\max_{i}\{1/\beta_{i},i=1,2,\dots,d\}.$
@@ 5.1. Data Partitioning and Scaling
7:  Dimension scaling, $\bm{X}_{p,j}:=\bm{X}^{org}_{p,j}/\beta_{j}$, $j=1,2,\ldots,d$.
@@
9:Return: Redistributed and scaled $\bm{X}_{1:P}$
5.2. Random Anchor Clustering (RAC)
Next, small blocks $\bm{B}_{p}$ are created, defined as $\bm{B}_{p}:=\{\bm{B}_{p,1},\ldots,\bm{B}_{p,k_{p}}\}$, for each partition by RAC independently. The RAC method is detailed in Algorithm 3. Here, we replace K-means clustering in block Vecchia GPs (Pan et al., 2025) with RAC to reduce the computational cost incurred by K-means for large datasets while maintaining comparable approximation accuracy. The RAC algorithm begins by randomly selecting $k_{p}$ anchors (centers) for each worker $p=1,2,\dots,P$ in parallel. These anchors serve as the centers of blocks. Subsequently, each $\bm{x}_{p,i}$ is assigned to the $j$th block based on their distance, as shown in line 5 of Algorithm 3, thus constructing clusters, $\bm{B}_{p}:=(\bm{B}_{p,1},\bm{B}_{p,2},\ldots,\bm{B}_{p,k_{p}})$ with $p=1,2,\dots,P$. The RAC method does not involve any communication between workers, considering that the size of clusters is typically much smaller than the overall dataset. For example, a single cluster may contain only 100 locations, whereas the entire dataset comprises 10 million locations. In Step 2 of Figure 2, each worker creates 4 blocks in their data partition.
Algorithm 3 RAC algorithm $\mathcal{C}(\bm{X}_{1:P})$
1:Input: $\bm{X}_{1:P}$ full dataset.
2:Output: $\bm{B}_{1:P}$ set of cluster for each node.